# WebXR, A-Frame and Networked-Aframe as a Basis for an Open Metaverse: A Conceptual Architecture

### v3

@@ Abstract
Metaverse, WebXR, Immersive experience, Spatial computing, Extended reality, Standards, World Wide Web, Browsers, Open source software.
@@ I Introduction
In the burgeoning realm of virtual reality and digital interconnectedness, the vision of an open metaverse—a vast, shared, and interoperable virtual space—has captivated the imagination of technologists, creators, and users alike. This work sets out to propose a WebXR-based cross-platform conceptual architecture designed to serve as a cornerstone for building such an open metaverse, delivered through the World Wide Web. Central to this exploration is the research question:
@@ II-B WebXR
A constantly updated list of browsers supporting WebXR is available online [^3] https://caniuse.com/webxr and, at the time of writing (2024), it shows that the latest versions of Chrome, Edge, Opera, and Samsung Internet, fully support the API. Firefox and Safari also support the API—not yet on iOS [^4] In the meantime, WebXR is supported by other iOS apps such as Mozilla’s WebXR Viewer.—although it may have to be manually enabled by the user. Consequently, all the most popular XR headsets also support WebXR, through their built-in browsers: Meta Quest, Microsoft HoloLens, Apple Vision Pro, HTC Vive, Samsung Gear, Google Cardboard and others. Interestingly, desktop and mobile browsers also support WebXR; however, since desktops, laptops, smartphones and smart TVs are not specifically designed for XR, the browser is forced to graphically simulate the immersive features of WebXR, depending on the screen and the platform. A-Frame apps can show a button in the lower right corner (Fig. 1)
@@ III Previous work
Researchers started to identify WebXR, A-Frame, and Networked-Aframe as promising technologies for the metaverse in 2023, albeit with limited exploration in scholarly articles so far. Earlier works about Networked-Aframe and virtual worlds such as Mozilla Hubs (without mentioning the metaverse) date back to 2019.
@@ V Conceptual architecture
The first three components are encapsulated within a browser environment due to the stack’s capability to deliver immersive experiences directly through a web browser without the need for external applications.
@@ VI Features of a WebXR-based metaverse
According to the definitions provided in section 1, persistence and multi-user concurrency are two key features of the metaverse. Besides, Mozilla’s engineers described several “open problems” of WebXR [18], which can be reformulated and classified into five groups: immersive web browsers, computer vision, geolocation, persistence, and multi-user concurrency. Let us see how these critical issues can be addressed and overcome in a state-of-the-art WebXR-based metaverse.
@@ VI-A Immersive browsers
In 2022, after WebXR was integrated into all the most popular web browsers, Mozilla ceased to maintain a browser specifically developed for VR devices, named Firefox Reality: an additional immersive browser is no longer considered essential because all the major browsers, on all the major operating systems, [^6] Except for Safari on iOS, as seen in section 4. are now able to provide immersive experiences, as indicated in the bottom layers of Fig. 3.
@@ VI-B Persistence
This storage modality operates on the back end, typically hosted in the cloud or on dedicated servers, facilitating centralized management of data that needs to be persistently available across sessions and platforms. Server-side databases, whether SQL-based (such as PostgreSQL or MySQL) for structured data or NoSQL databases (like MongoDB, Couchbase, or DynamoDB) for unstructured or semi-structured data, offer scalable solutions for managing complex datasets inherent in virtual worlds.
@@ VI-C Multi-user concurrency
It supports the representation of users within the virtual space, often through avatars, allowing for the visualization of other users’ presence and movements. This feature is crucial for social interactions and collaborative activities within the metaverse.
@@ VII Implementation and prototyping
Following the conceptual architectural for spatial web apps, illustrated in Fig. 3, a prototype was developed to demonstrate the practical application of the proposed stack. The prototype aimed to showcase a basic yet functional open metaverse environment, focusing on aspects such as user interaction, networked multiplayer capabilities, and cross-platform accessibility. The implementation phase of the research focused on leveraging Mozilla Hubs [18], an open-source project available on GitHub, as a prototype to demonstrate the practical application of the conceptual architecture for an open metaverse. Mozilla Hubs represents a compelling example of how WebXR, A-Frame, and Networked-Aframe can be integrated to create immersive, cross-platform virtual environments. Through the implementation and prototyping phase, the goal was to demonstrate not only the feasibility of the conceptual architecture but also its practical value in creating accessible, immersive virtual environments. The implementation process encompassed the following steps:
VII-A Selection of Mozilla Hubs
Mozilla Hubs was chosen due to its alignment with the architectural vision, leveraging WebXR for virtual reality experiences directly in the browser, A-Frame for creating 3D scenes with HTML, and Networked-Aframe for enabling real-time, multiplayer interactions. Its open-source nature allowed for in-depth analysis and experimentation, providing a solid foundation for the prototype.
@@ VII-B Customization and deployment
To test the conceptual architecture, a Mozilla Hubs environment was customized to suit the research objective. This involved modifying existing scenes to incorporate unique elements that would test the limits of the architecture, such as complex geometries and high-density multiplayer interactions. Following customization, the environment was deployed on a server to facilitate access and interaction from various devices and platforms.
@@ VIII-A Technical performance metrics
The technical performance of the Mozilla Hubs prototype was evaluated using several key metrics:
@@ VIII-B User experience survey
To gather qualitative feedback on the user experience, participants were invited to explore the Mozilla Hubs environment and engage in various activities. The results were as follows:
@@ IX Discussion
The evaluation of the Mozilla Hubs prototype, grounded in the proposed conceptual architecture for an open metaverse, provided significant insights into the practical application and potential of WebXR, A-Frame, and Networked-Aframe. This discussion elaborates on the findings, addressing the strengths, limitations, and implications for future research and development.
@@ X Conclusion
This work proposed a WebXR-based cross-platform conceptual architecture for developing spatial web apps using the A-Frame and Networked-Aframe frameworks with a view to an open, accessible, and interoperable metaverse. A Mozilla Hubs prototype was implemented and evaluated, supporting the capability of the proposed technology stack to enable immersive experiences across different platforms and devices. Positive feedback on ease of use of the immersive environment further corroborates the proposed approach, underscoring its effectiveness in facilitating engaging and interactive virtual spaces.

### v5

@@ Abstract
Metaverse, WebXR, Immersive experience, Spatial computing, Extended reality, Open standards, World Wide Web, Browsers.
@@ I Introduction
In the growing area of virtual reality and digital interconnectedness, the vision of an open metaverse—a vast, shared, and interoperable virtual space—has captivated the imagination of technologists, creators, and users alike. This work sets out to propose a WebXR-based cross-platform conceptual architecture designed to serve as a cornerstone for building such an open metaverse, delivered through the World Wide Web. Central to this exploration is the research question:
@@ II-B WebXR
A constantly updated list of browsers supporting WebXR is available online [^3] https://caniuse.com/webxr and, at the time of writing (2024), it shows that the latest versions of Chrome, Edge, Opera, and Samsung Internet, fully support the API. Firefox and Safari also support the API—not yet on iOS [^4] In the meantime, WebXR is supported by other iOS apps such as Mozilla’s WebXR Viewer.—although it may have to be manually enabled by the user. Consequently, all the most popular XR headsets also support WebXR, through their built-in browsers: Meta Quest, Microsoft HoloLens, Apple Vision Pro, HTC Vive, Samsung Gear, Google Cardboard and others. Interestingly, desktop and mobile browsers also support WebXR; however, since desktops, laptops, smartphones and smart TVs are not specifically designed for XR, the browser is forced to graphically simulate the immersive features of WebXR, depending on the screen and the platform. This is why A-Frame [^5] See section II-C. apps can show a button in the lower right corner (fig. 1)
@@ III Previous work
Researchers started to identify WebXR, A-Frame, and Networked-Aframe as promising technologies for the metaverse in 2023, albeit with limited exploration in scholarly articles so far. Earlier works about Networked-Aframe and virtual worlds such as […] (without mentioning the metaverse) date back to 2019.
@@ V Conceptual architecture
As seen in section II-B, the first three components are natively supported by web browsers; therefore, in fig. 3, they are encapsulated within a “browser” environment, indicating the stack’s capability to deliver immersive experiences directly through a web browser without the need for external applications.
@@ VI Features of a WebXR-based metaverse
According to the definitions provided in section II-A, persistence and multi-user concurrency are two key features of the metaverse. Besides, […]’s engineers described several “open problems” of WebXR […], which can be reformulated and classified into five groups: immersive web browsers, computer vision, geolocation, persistence, and multi-user concurrency. Let us see how these critical issues can be addressed and overcome in a state-of-the-art WebXR-based metaverse.
@@ VI-A Immersive browsers
In 2022, after WebXR was integrated into all the most popular web browsers, Mozilla ceased to maintain a browser specifically developed for VR devices, named Firefox Reality: an additional immersive browser is no longer considered essential because all the major browsers, on all the major operating systems, [^7] Except for Safari on iOS, as seen in section II-B. are now able to provide immersive experiences, as indicated in the bottom layers of fig. 3.
@@ VI-B Persistence
This storage modality operates on the back end, typically hosted in the cloud or on dedicated servers, facilitating centralized management of data that needs to be persistently available across sessions and platforms. Server-side databases, whether SQL-based (such as PostgreSQL or MySQL) for structured data or NoSQL databases (such as MongoDB, Couchbase, or DynamoDB) for unstructured or semi-structured data, offer scalable solutions for managing complex datasets inherent in virtual worlds.
@@ VI-C Multi-user concurrency
Networked-Aframe supports the representation of users within the virtual space, often through avatars, allowing for the visualization of other users’ presence and movements. This feature is crucial for social interactions and collaborative activities within the metaverse.
@@ VII Implementation and prototyping
Following the conceptual architectural for spatial web apps, illustrated in fig. 3, a prototype was developed to demonstrate the practical application of the proposed stack. The prototype aimed to showcase a basic yet functional open metaverse environment, focusing on aspects such as user interaction, networked multiplayer capabilities, and cross-platform accessibility. The implementation phase of the research focused on leveraging […], an open-source project available on GitHub, as a prototype to demonstrate the practical application of the conceptual architecture for an open metaverse. […] represents a compelling example of how WebXR, A-Frame, and Networked-Aframe can be integrated to create immersive, cross-platform virtual environments. Through the implementation and prototyping phase, the goal was to demonstrate not only the feasibility of the conceptual architecture but also its practical value in creating accessible, immersive virtual environments. The implementation process encompassed the following steps:
VII-A Selection of […]
[…] was chosen due to its alignment with the architectural vision, leveraging WebXR for virtual reality experiences directly in the browser, A-Frame for creating 3D scenes with HTML, and Networked-Aframe for enabling real-time, multiplayer interactions. Its open-source nature allowed for in-depth analysis and experimentation, providing a solid foundation for the prototype.
@@ VII-B Customization and deployment
To test the conceptual architecture, a […] environment was customized to suit the research objective. This involved modifying existing scenes to incorporate unique elements that would test the limits of the architecture, such as complex geometries and high-density multiplayer interactions. Following customization, the environment was deployed on a server to facilitate access and interaction from various devices and platforms.
@@ VIII-A Technical performance metrics
The technical performance of the […] prototype was evaluated using several key metrics:
@@ VIII-B User experience survey
To gather qualitative feedback on the user experience, participants were invited to explore the […] environment and engage in various activities. The results were as follows:
@@ IX Discussion
The evaluation of the […] prototype, grounded in the proposed conceptual architecture for an open metaverse, provided significant insights into the practical application and potential of WebXR, A-Frame, and Networked-Aframe. This discussion elaborates on the findings, addressing the strengths, limitations, and implications for future research and development.
@@ X Conclusion
This work proposed a WebXR-based cross-platform conceptual architecture for developing spatial web apps using the A-Frame and Networked-Aframe frameworks with a view to an open, accessible, and interoperable metaverse. A […] prototype was implemented and evaluated, supporting the capability of the proposed technology stack to enable immersive experiences across different platforms and devices. Positive feedback on ease of use of the immersive environment further corroborates the proposed approach, underscoring its effectiveness in facilitating engaging and interactive virtual spaces.